How much 8% saline solution should Kent mix with 80 cubic centimeters of an 18% saline solution to produce a 12% saline solution?
Answer by mananth(16946)

------ Percent ----------------quantity
Saline solution I 18 ----------------80
Saline solution II 8 ----------------x
Total 12 ----------------80+x
18*80+8*x=12(80 +x)
1440+8x =960+12x
8x-12x=960-1440

-4x=-480
/-4
x=120 cu.cm 8% Saline solution II